### v1

# SAEMark: Multi-bit LLM Watermarking with Inference-Time Scaling

Abstract
Watermarking LLM-generated text is critical for content attribution and misinformation prevention. However, existing methods compromise text quality, require white-box model access and logit manipulation—limitations that exclude API-based models and multilingual scenarios. We propose SAEMark, a general framework for post-hoc multi-bit watermarking that embeds personalized messages solely via inference-time, feature-based rejection sampling without altering model logits or requiring training. Our approach operates on deterministic features extracted from generated text, selecting outputs whose feature statistics align with key-derived targets. This framework naturally generalizes across languages and domains while preserving text quality through sampling LLM outputs instead of modifying. We provide theoretical guarantees relating watermark success probability and compute budget that hold for any suitable feature extractor; empirically, we demonstrate the framework’s effectiveness using Sparse Autoencoders (SAEs), achieving superior detection accuracy and text quality. Experiments across 4 datasets show SAEMark’s consistent performance, with 99.7% F1 on English and strong multi-bit detection accuracy. SAEMark establishes a new paradigm for scalable watermarking that works out-of-the-box with closed-source LLMs while enabling content attribution.  [^1] We open-source code and data at: https://zhuohaoyu.github.io/SAEMark
1 Introduction
Large language models (LLMs) have revolutionized text generation across domains, from creative writing to code synthesis (Brown et al., 2020; Guo et al., 2024). However, their ability to produce human-quality text at scale raises serious concerns about misinformation, copyright infringement, and content laundering. As these models become ubiquitous, reliably attributing AI-generated content becomes critical for accountability and trust.
Watermarking—embedding detectable signatures into generated text—offers a promising solution, but existing approaches face a fundamental tradeoff. They must preserve text quality while enabling reliable detection, operate across languages and domains, and scale to distinguish between many users or sources. Most critically, they must work with real-world deployment constraints where model providers offer only API access without exposing internal parameters.
The challenge becomes even more complex for multi-bit watermarking. Beyond simply detecting AI-generated text, the goal is to encode and recover a specific message $m\in\{0,1\}^{b}$—such as a user identifier for personalized attribution. This enables answering not just “is this AI-generated?” but “which specific user or system generated this text?” Such fine-grained attribution is essential for large-scale deployment where accountability matters.
Existing watermarking methods struggle with these requirements. Token-level approaches like KGW (Kirchenbauer et al., 2023) and EXP (Aaronson & Kirchner, 2022) require direct access to model logits, excluding API-based deployment, and can degrade text quality through probability manipulation. Syntactic methods (Hou et al., 2023) fail to generalize across languages, while specialized approaches (Lee et al., 2024) work well in narrow domains but break down when applied more broadly. Even recent black-box methods (Bahri & Wieting, 2024; Chang et al., 2024) rely on surface-level statistics or require auxiliary models, limiting their robustness and scalability.
We introduce SAEMark, a fundamentally different approach that sidesteps these limitations entirely. Our key insight is deceptively simple: different LLM generations exhibit distinct patterns in their semantic features, and these patterns can be leveraged for watermarking through selection rather than modification. Instead of altering how text is generated, we generate multiple candidates and choose those whose feature patterns align with a watermark key.
This approach works by operating on meaningful units of text—sentences for natural language, functions for code. For each unit, we extract deterministic features that capture semantic properties, compute a scalar statistic, and normalize it to behave predictably across different texts. Using the watermark key, we derive target values for each position. During generation, we sample multiple candidates from the LLM and select the one whose feature statistic is closest to the target, ensuring the final sequence encodes the desired message.
The elegance lies in what we don’t change: no model weights, no logit manipulation, no token modifications. Every selected text segment is a natural LLM output, preserving quality while enabling attribution. The approach works with any LLM through API calls, generalizes across languages and domains, and provides theoretical guarantees on watermark success that scale predictably with computational budget.
Our contributions span theory and practice. We develop a general framework for watermarking through feature-guided selection that works with any feature extractor and any language model API. We provide theoretical guarantees that explain how watermark success scales with computational resources and text length, independent of the specific features used. Finally, we demonstrate a practical instantiation using Sparse Autoencoders that achieves superior detection accuracy and text quality across English, Chinese, and code, encoding more information per unit length than existing multi-bit approaches.
Figure 1: An overview of SAEMark.
2 Preliminaries
2.1 Related Work
LLM watermarking is a technique to embed special patterns into the output of LLMs, and has traditionally been used to identify LLM generated text from human-written text (Jawahar et al., 2020). Different from post-hoc detection methods (Zellers et al., 2019) that analyze statistical patterns in existing text, language model watermarking aims to embed detectable signatures during generation (Kirchenbauer et al., 2023).
Existing approaches exhibit several limitations. Many methods compromise generation quality through direct manipulation of token probabilities (Kirchenbauer et al., 2023) or syntactic modifications (Atallah et al., 2002). The challenge of language and domain generalization remains largely unaddressed, with current techniques primarily optimized for English and struggling with multilingual content or specialized domains like code (Lee et al., 2024). Notably, PersonaMark (Zhang et al., 2024b) represents early attempts at personalized watermarking, but its reliance on English-specific syntactic patterns and closed-source implementation makes scalability and cross-lingual capability difficult to verify. Recently, more multi-bit watermarking methods have been proposed to embed multiple bits of information into generated text (Lau et al., 2024; Zhang et al., 2024a; Yoo et al., 2023a; Guan et al., 2024; Yoo et al., 2023b; Qu et al., 2024), primarily by extending single-bit watermarking that manipulates logits during generation; these methods inherit the limitations of single-bit designs.
Complementary black-box watermarks avoid white-box logit access by using post-hoc selection or rewriting (Chang et al., 2024). However, they typically operate on surface statistics or introduce auxiliary model dependencies and do not directly address multi-bit message embedding at scale. Our framework differs by performing inference-time selection among naturally generated candidates using deterministic feature statistics computed over domain-appropriate units. This enables extractor-agnostic analysis and multilingual, domain-agnostic multi-bit watermarking without modifying model logits.
2.2 Sparse Autoencoders
Sparse Autoencoders (SAEs) are pre-trained interpretability tools that decompose LLM activations into human-understandable features (Bricken et al., 2023). For a given base model $\mathcal{M}$ and layer $l$, an SAE processes hidden states $\mathbf{h}_{t}$ at position $t$ as:
$\mathbf{f}_{t}=\text{SAE}_{l}(\mathbf{h}_{t})$ (1)
where $\mathbf{f}_{t}\in\mathbb{R}^{m}$ is a sparse vector (typically $m\gg\dim(\mathbf{h}_{t})$) with $\leq$ 5% active features. The SAE is trained through two objectives: 1) reconstruct original activations, and 2) enforce feature sparsity via $L_{1}$ regularization:
$\mathcal{L}=\underbrace{\|\mathbf{h}_{t}-\text{Dec}(\mathbf{f}_{t})\|^{2}}_{L_{rec}}+\lambda\underbrace{\|\mathbf{f}_{t}\|_{1}}_{L_{sparse}}$ (2)
This training produces features that correspond to interpretable concepts (Bricken et al., 2023; Templeton et al., 2024). For instance, SAEs applied to Gemma 2B (Team et al., 2024)’s final transformer layer, GemmaScope, reveal features like ”Python function definitions” or ”Concept related to color blue” (Lieberum et al., 2024).
Our watermarking leverages three key properties of pre-trained SAE features. First, layer-specific patterns capture distinct behaviors from different model layers. Second, multilingual activation allows the same features to fire for equivalent concepts across languages. Third, sparsity enables efficient analysis through few active features per token. These properties support language-agnostic statistics via masked feature aggregations:
$\phi(\mathbf{y})=\frac{1}{|\mathbf{y}|}\sum_{t}\mathbf{f}_{t}\odot\mathbf{m}$ (3)
where $\mathbf{m}$ filters background features that fire ubiquitously regardless of content (e.g., punctuation-associated features). The summary $\phi(\mathbf{y})$ provides a deterministic statistic used by our generation and detection procedures (Sec 3).
2.3 Task Definition
We adopt a multi-bit view of attribution: beyond binary detection, the objective is to encode a message $m\in\{0,1\}^{b}$ that is recoverable at detection. Personalized attribution is a special case where $m$ encodes a user identifier bound to a key.
Generation (multi-bit).
Given a base LLM $\mathcal{M}$, watermark key $k\in\mathcal{K}$, input $\mathbf{x}$, and message $m\in\{0,1\}^{b}$, the algorithm produces $\mathbf{y}$ by post-hoc selection over $\mathcal{M}$’s outputs (black-box/API-compatible; no access to logits or parameters; cf. black-box watermarking (Bahri & Wieting, 2024; Chang et al., 2024)):
$\mathbf{y}=\text{Mark}(\mathcal{M},\mathbf{x},k,m).$ (4)
Detection (multi-bit).
For any text $\mathbf{y}^{\prime}$ and key $k$, the detection algorithm outputs a decoded message or reject:
[…]
The scheme targets three properties: key privacy (deriving $k$ from watermarked outputs is hard), verifier-held detectability (any party holding $k$ can verify), and collusion resistance (multiple keys should not facilitate removal or forgery).
Threat model and scope.
Our focus is attribution without storing sensitive content. This work does not claim cryptographic unforgeability when keys are known; preventing adversarial forgeries under black-box constraints is an important direction for security-focused follow-ups.
3 Methodology
We present a general framework for post-hoc, multi-bit watermarking via feature-based rejection sampling. The key observation is that different LLM generations produce distinct values of deterministic feature statistics computed over domain-appropriate units, and these statistics can be steered by selecting among naturally generated candidates, without modifying model logits, parameters or generated texts. We structure the section as follows: (1) a general framework that is extractor-agnostic, (2) theoretical guarantees with an emphasis on worst-case bounds, and (3) an effective instantiation using sparse autoencoders as feature extractors.
3.1 General Framework for Feature-based Watermarking
Our approach operates on a simple intuition: suppose we have a deterministic feature extractor that maps any text sequence into a scalar value, where such values follow a predictable distribution (e.g., approximately normal) for naturally generated text. Given a watermark key $k$ encoding multi-bit information, we can derive a sequence of target scalar values from this key. During generation, we produce text chunk by chunk, ensuring each chunk yields a scalar value with the smallest difference to its corresponding target—effectively implementing rejection sampling guided by our feature-based reward function. This process steers generation toward key-dependent patterns without modifying the underlying language model.
Text segmentation.
We segment text into smaller units $\{u_{i}\}_{i=1}^{M}$ such as sentences for natural language or function blocks for code. Each unit will carry one symbol of the watermark signal.
Feature extraction.
A deterministic feature extractor $\phi:\mathcal{U}\to\mathbb{R}^{d}$ maps each text unit to a feature vector, from which we compute a scalar statistic $s(u)=g(\phi(u))\in\mathbb{R}$. Crucially, we assume that this statistic follows a predictable distribution when computed over naturally generated text units.
Statistical normalization.
To enable analysis independent of the specific feature extractor, we normalize the statistic $s(u)$ to a standard range $[0,1]$ using its empirical distribution. Specifically, we estimate the cumulative distribution function $F_{S}$ from natural text, then map each unit’s statistic via $z(u)=\hat{F}(s(u))$ where $\hat{F}$ is the empirical CDF. This ensures $z(u)$ values are approximately uniformly distributed for natural text.
Watermark generation process.
Given a watermark key $k$ encoding multi-bit information, we first randomly generate a sequence of target values $\{\tau_{i}\}_{i=1}^{M}$ by seeding a PRNG generator with $k$ and sampling each $\tau_{i}$ from a suitable range deterministically. Then, for each position $i$, we generate $N$ candidate text units from the LLM and select the candidate $c^{*}$ whose normalized statistic $z(c^{*})$ is closest to the target $\tau_{i}$.
Watermark detection process.
To detect a watermark in input text, we segment it into units, compute the normalized statistic $z(u)$ for each unit, and compare the resulting sequence $\{z_{i}\}$ against target sequences derived from candidate keys. We apply a two-stage CheckAlignment process to verify sequence before statistical testing.
The CheckAlignment process employs two critical filters to ensure the observed sequence $\{z_{i}\}_{i=1}^{M}$ and expected target sequence $\{\tau_{i}\}_{i=1}^{M}$ are sufficiently similar:
Range Similarity Filter: This constraint ensures the dynamic ranges of observed and target sequences are similar:
$R_{\min}<\frac{z_{\max}-z_{\min}}{\tau_{\max}-\tau_{\min}}<R_{\max}$ (6)
where $z_{\max}=\max_{i}z_{i}$, $z_{\min}=\min_{i}z_{i}$, and similarly for $\tau$. We typically set $R_{\min}=0.95$, $R_{\max}=1.05$.
Overlap Rate Filter: This constraint ensures sufficient overlap between the value ranges of both sequences:
$\frac{|\{i:\tau_{i}\in[z_{\min},z_{\max}]\}|}{M}\geq O_{\min}$ (7)
where $M$ denotes the number of textual units in the sequence and $O_{\min}=0.95$ ensures that at least 95% of target values fall within the observed range.
These two filters aim to eliminate spurious matches: the range similarity filter prevents matching sequences with fundamentally different statistical properties, while the overlap rate filter ensures meaningful correspondence between target and observed values. Only after passing both alignment checks do we apply Student’s t-test for statistical significance. The key with the highest significance score is returned if it passes the threshold; otherwise, we classify the text as unwatermarked.
Input: Prompt $c$, key $k$, LLM $G$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, units $M$, attempts $K$, candidates NN, alignment thresholds
Output: Watermarked text $x^{*}$
for $attempt\leftarrow 1$ to $K$ do
 $x^{*}\leftarrow c$; $\{\tau_{i}\}_{i=1}^{M}\leftarrow\text{TargetsFromKey}(k)$; $\{z_{i}\}\leftarrow\emptyset$
 for $i\leftarrow 1$ to $M$ do
    $\mathcal{X}\leftarrow\text{GenerateCandidates}(G,x^{*},N)$
    $x_{best}\leftarrow\arg\min\limits_{x\in\mathcal{X}}\big{|}\hat{F}(s(\phi(x))){-}\tau_{i}\big{|}$
    $x^{*}\leftarrow x^{*}\oplus x_{best}$; $z_{i}\leftarrow\hat{F}(s(\phi(x_{best})))$
   end for
 if CheckAlignment($\{\tau_{i}\},\{z_{i}\}$then
    return $x^{*}$
   end if
end for
return $x^{*}$
Algorithm 1 Watermark generation
Input: Text $x$, candidate keys $\mathcal{K}$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, alignment thresholds, significance level $\alpha$
Output: Detection result $d\in\mathcal{K}\cup\{\emptyset\}$
$\{z_{j}\}\leftarrow[\hat{F}(s(\phi(u)))\;\forall u\in\text{SegmentByDomain}(x)]$
$\mathcal{D}\leftarrow\emptyset$
foreach $k_{i}\in\mathcal{K}$ do
 $\{\tau_{j}\}\leftarrow\text{TargetsFromKey}(k_{i},|\{z_{j}\}|)$
 if CheckAlignment($\{\tau_{j}\},\{z_{j}\}$then
    $t,p\leftarrow\text{StudentTTest}(\{z_{j}\},\{\tau_{j}\})$
    if $t>t_{\alpha/2}\land p<\alpha$ then
       $\mathcal{D}\leftarrow\mathcal{D}\cup\{(k_{i},t)\}$
      end if
   end if
end foreach
return $\operatorname{argmax}_{(k_{i},t_{i})\in\mathcal{D}}t_{i}$ if $\mathcal{D}\neq\emptyset$ else $\emptyset$
Algorithm 2 Watermark detection
Figure 2: Pseudocode for SAEMark: generation and detection.
3.2 Theoretical Analysis and Guarantees
We provide theoretical guarantees on watermark embedding success that enable reliable detection by a conservative bound. For clarity, we present our analysis for a single textual unit and refer to our experiments for empirical validation of multi-unit performance with CheckAlignment process.
Embedding success under Gaussian assumption.
Let target values $\tau$ be sampled from the feasible range $\mu-2\sigma,\mu+2\sigma]$ where the feature statistic follows $S\sim\mathcal{N}(\mu,\sigma^{2})$. Given $N$ candidate generations with feature statistics $S_{1},S_{2},\ldots,S_{N}$, we seek the probability of finding at least one candidate within relative tolerance $k$ of our target:
$\mathbb{P}(\exists j:|S_{j}-\tau|\leq k\tau)\geq 1-(1-p_{\min})^{N}$ (8)
Worst-case analysis and tight bounds.
To derive conservative guarantees, consider the worst-case target $\tau=\mu+2\sigma$ at the boundary of the feasible range. The single-candidate success probability becomes:
$p_{\min}=\mathbb{P}((1-k)\tau\leq S_{j}\leq(1+k)\tau)=\Phi\left(\frac{(1+k)(\mu+2\sigma)-\mu}{\sigma}\right)-\Phi\left(\frac{(1-k)(\mu+2\sigma)-\mu}{\sigma}\right)$ (9)
where $\Phi$ denotes the standard normal CDF. This simplifies to $p_{\min}=\Phi(2(1+k)+k\mu/\sigma)-\Phi(2(1-k)-k\mu/\sigma)$.
The fundamental insight is that observed feature statistics are tightly bounded to target values. Setting strict tolerance $k$ guarantees strong detection accuracy: embedding succeeds with high probability $1-(1-p_{\min})^{N}$ even with conservative parameters, while detection maintains precision because legitimate watermarks exhibit tight statistical binding that unwatermarked text cannot match. This framework provides exponential improvement with candidate count $N$, enabling principled compute-accuracy tradeoffs validated empirically across diverse tasks.
3.3 Sparse Autoencoder Instantiation
What concrete feature extractor should we use? We need statistics that are deterministic, semantically meaningful, and statistically regular. Sparse autoencoders—interpretability tools designed to understand language model internals—provide an ideal solution. They decompose language representations into interpretable semantic components (”technical writing,” ”narrative voice,” ”mathematical reasoning”) that exhibit distinctly different activation patterns across generations. By applying the sparse autoencoder to a separate ”anchor” model, our approach remains compatible with any target language model, including API services, while extracting the discriminative yet predictable statistics our framework requires.
Figure 3: Distribution analysis of FCS. FCS distribution with density estimation (left) and Q-Q plot (right); statistical tests support approximate normality.
The Feature Concentration Score intuition.
Rather than using raw sparse autoencoder outputs, we compute a Feature Concentration Score (FCS) that captures a fundamental property of coherent text: semantic focus. The key insight is that well-formed text tends to concentrate its semantic activation on a consistent set of relevant features, while unfocused or incoherent text spreads activation more uniformly. For example, a technical manual concentrates activation on features related to formal language and domain expertise, while creative writing focuses on narrative and stylistic features.
This concentration pattern provides an ideal watermark signal—we can steer generation toward specific concentration levels without affecting text quality, since both high and low concentration can correspond to natural, well-written text in different contexts. The FCS measures this by identifying the most salient feature activated by each token, then computing what fraction of the total activation mass concentrates in these important features:
$\text{FCS}(T)=\frac{\sum_{t=1}^{n}\sum_{i\in S}\phi_{t,i}}{\sum_{t=1}^{n}\|\phi_{t}\|_{1}},$ (10)
where $S=\{\arg\max_{i}(\phi_{t,i}\odot m_{i}):t=1,\ldots,n\}$ contains the indices of the most salient features across all tokens, after applying the background mask $m$ and deduplication. This provides our framework’s statistic $s(u)=\text{FCS}(u)$, which empirically follows approximately normal distributions across different domains and languages, validating our theoretical assumptions. We provide illustration and detailed analysis of this process in Appendix D.
Implementation details.
Two practical optimizations bridge the gap between our theoretical framework and robust empirical performance. First, the CheckAlignment algorithm’s eliminate spurious statistical matches that would otherwise compromise detection accuracy. Second, background feature masking ensures FCS calculations focus on discriminative semantic patterns rather than ubiquitous surface features. We precompute a mask excluding ”background” SAE features, like those related to punctuation or basic grammar, to focus on discriminative semantic patterns. With empirically observed parameters $\mu=0.142$, $\sigma=0.029$, and tolerance $\epsilon=0.1$, our bounds yield concrete success probabilities: $N=50$ achieves $99\%$ per-unit success, $N=20$ maintains $85.32\%$, and $N=10$ achieves $61\%$. Since each generation involves multiple units, overall success rates significantly exceed these per-unit bounds. Modern inference engines support parallel generation of $N$ candidates simultaneously, making the approach practically efficient despite these extra compute overhead. We have extensive ablations and empirical results in the following experiments.
4 Experiments
Our experiments systematically address four fundamental questions: (1) How accurate and quality-preserving is our method compared to existing single-bit and multi-bit watermarks? (2) What are the computational overhead characteristics and scalability properties in practice? (3) How robust is our method against adversarial attacks? (4) Which components contribute most significantly to bridging the gap between theoretical bounds and empirical performance?
4.1 Experimental Setup
Table 1: Dataset Statistics. Characteristics of the multilingual benchmarks used in evaluation.
C4 (2020) LCSTS (2015) MBPP (2021) PandaLM (2023b)
# Samples 500 500 257 169
Language English Chinese Python English
Task Type
Completion
Summarization
Code Gen.
Inst. Following
From test split of sanitized version of MBPP.
Setup.
We evaluate on 4 diverse datasets as shown in Table 1. Following common practice in prior work, we report Accuracy, Recall, F1 at 1% FPR. For text quality, we report win-rates of pairwise comparison on PandaLM judged by GPT-4o in our main results, and average pointwise scores on BIGGen-Bench judged by their officially released judge model as an alternative text-quality experiment. We use implementation for single-bit watermarks from MarkLLM (Pan et al., 2024) toolkit and Waterfall (Lau et al., 2024) as it’s the current best open-source training-free multi-bit watermark similar to our setting. Full details of baselines in Appendix C.
4.2 Watermarking Accuracy and Text Quality
Multi-bit watermarking poses a fundamentally harder challenge than single-bit detection: we must embed significantly more information into the same text length while maintaining both accuracy and quality. Despite this increased difficulty, Table 2 shows SAEMark achieves superior accuracy compared to both single-bit baselines and the current best multi-bit watermark across all domains.
Table 2: Comparison of Watermarks. We generate watermarked and unwatermarked texts and then report detection performance at 1% false positive rate (FPR), all in single-bit settings. Best results are in bold and second-best are underlined. All metrics are reported as percentages (%).
Method C4 (English, 2020) LCSTS (Chinese, 2015) MBPP (Code, 2021) PandaLM (Instruction,  2023b)
[…]
Single-bit Watermarks
KGW (2023) 99.2 99.6 99.2 99.1 98.8 99.1 65.4 31.9 48.0 41.5 89.9 80.4 88.8
EXP (2022) 99.5 99.6 99.5 99.3 99.4 99.3 57.8 16.7 28.4 23.2 79.3 59.4 74.2
UPV (2023) 86.0 72.0 83.7 90.5 91.0 90.5 51.6 3.1 6.0 36.0 54.0 8.0 14.8
Unigram (2023) 98.8 98.6 98.8 98.2 97.0 98.2 65.4 31.9 48.0 35.3 53.3 7.2 13.4
DIP (2023) 96.0 92.6 95.9 97.7 96.2 97.7 60.7 22.6 36.5 36.5 81.5 63.8 77.5
Unbiased (2023b) 96.7 94.4 96.6 97.8 96.4 97.8 64.0 29.2 44.8 40.2 74.3 49.3 65.7
SynthID (2024) 98.2 97.2 98.2 97.6 96.2 97.6 62.5 26.1 41.0 36.0 81.2 63.0 77.0
SWEET (2024) 99.6 99.6 99.6 50.0 0.0 0.0 72.4 45.9 62.4 47.2 87.7 76.8 86.2
Multi-bit Watermarks
Waterfall (2024) 93.6 88.0 93.2 95.3 91.6 95.1 52.5 6.2 11.6 46.4 73.2 47.1 63.7
SAEMark (Ours) 99.7 99.8 99.7 99.2 99.6 99.2 74.5 50.2 66.3 67.6 86.6 73.9 84.6
Accuracy across domains.
SAEMark establishes new state-of-the-art performance: 99.7% F1 on English, 99.2% on Chinese, and 66.3% on code. Notably, we outperform specialized methods in their own domains—surpassing code-specific SWEET by 3.9 points F1 (66.3% vs. 62.4%) despite our general-purpose design. While other methods suffer severe cross-domain performance cliffs, SAEMark captures language-agnostic patterns that generalize across syntactic variations. The multi-bit comparison reveals particularly dramatic advantages: SAEMark outperforms the current best multi-bit method Waterfall by 6.5 points F1 on English (99.7% vs. 93.2%) and an exceptional 54.7 points on code (66.3% vs. 11.6%), demonstrating semantic feature-based selection’s clear superiority over vocabulary permutation approaches, especially in low-entropy domains.
Text quality.
Beyond accuracy, Table 2 shows SAEMark achieves the highest quality score (67.6%) on PandaLM as judged by GPT-4o pairwise comparisons. To study how this conclusion generalizes across different backbone models, since backbone performance is the most significant factor affecting text quality, we conduct additional evaluation on BIGGen-Bench comparing against both watermarked baselines and unwatermarked text.
Table 3: Text quality evaluation on BIGGen-Bench. Scores are on a 5-point Likert scale (higher is better), judged by the officially released BIGGen-Bench judge model (Kim et al., 2024).
Model Unwatermarked SAEMark KGW Waterfall
Qwen2.5-7B-Instruct 4.13 4.05 3.97 4.02
Llama-3.2-3B-Instruct 3.69 3.85 3.56 3.62
gemma-3-4b-it 4.26 4.23 3.98 4.19
Table 3 confirms SAEMark consistently achieves the highest quality among watermarking methods across three backbone models. This quality advantage stems from a fundamental difference in approach: rather than manipulating logits or applying external rewriting to obtain watermarked text, we simply select among naturally generated candidates. This post-hoc selection guarantees that every watermarked segment remains an unmodified LLM generation from the backbone model itself, ensuring text quality stays bounded by the its own capabilities.
4.3 Computational Overhead and Scalability
Our theoretical analysis suggested requiring N=50 candidates to achieve 99%+ accuracy per unit. However, through the two practical optimizations in our framework: background feature masking and CheckAlignment filters, we achieve strong performance with significantly reduced computational overhead in practice.
(a) Perf. vs. Sampled Candidates
N=5 N=10 N=20 N=50
C4 (Raffel et al., 2020)
Acc. 98.7 99.2 98.7 99.7
Rec. 77.4 96.8 98.7 99.8
F1 86.8 98.0 98.7 99.7
LCSTS (Hu et al., 2015)
Acc. 98.6 99.0 98.6 99.2
Rec. 72.6 96.0 98.0 99.6
F1 83.6 97.5 98.6 99.2
(b) Perf. vs. Average Latency
Method Acc. Rec. F1 Latency
KGW 99.0 99.5 98.9 3.24x
UPV 90.3 86.3 89.5 2.35x
DIP 99.5 99.7 99.5 3.29x
Waterfall 98.8 97.3 98.1 1.06x
Ours(N=50) 99.5 99.7 99.5 1.00x
Figure 4: Computational overhead analysis. (a) Performance vs. number of sampled candidates for SAEMark. (b) Performance vs. avg latency across different watermarks.
Practical efficiency.
Figure 4 (a) demonstrates this efficiency gain: N=10 achieves 98.0% F1 on English with reasonable overhead, while even N=5 attains 86.8% F1—substantially better than our conservative theoretical bounds predicted. This flexibility enables deployment across different computational budgets.
Moreover, subfigure (b) reveals a remarkable result: SAEMark achieves 99.5% F1 at 1.00× baseline latency, substantially outperforming methods requiring 3.24× latency (KGW) and 3.29× latency (DIP) for comparable accuracy.
Infrastructure advantage.
This performance difference reflects a genuine architectural advantage. Since SAEMark requires no logit manipulation, we can leverage highly optimized inference backends like TGI with parallel candidate generation and tricks like prefix caching and custom, optimized CUDA kernels. In contrast, these optimized frameworks do not provide efficient watermark implementations for logit-manipulation methods, as such implementations require significant backend rewriting and may impact performance. While this creates a significant latency difference despite some methods theoretically needing less compute overhead, we consider this a practical advantage reflecting the current state of inference infrastructure. Transparency note: We are being honest about these numbers—they reflect real deployment advantages rather than experimental artifacts.
Multi-bit scaling.
Figure 6 shows our approach maintains over 90% accuracy up to 10 bits (effectively differentiating 1,024 users) and 75% accuracy at 13 bits (8,192 users), substantially exceeding Waterfall’s performance through our high-dimensional SAE feature space. Importantly, this does not mean our method is only effective with 1,024 users—we are conducting fixed text length comparisons for fair evaluation. The superior information density stems from finer-grained semantic distinctions our framework enables.
4.4 Adversarial Robustness
Semantic SAE features provide inherent robustness against paraphrasing attacks. Figure 6 demonstrates our method’s resilience across three attack types—word deletion, synonym substitution, and context-aware substitution. SAEMark shows strong resilience to such attacks. Due to space limitations, extended results testing attack intensities up to 50% are provided in Appendix E, demonstrating continued robustness even under stronger attacks.
4.5 Ablation Studies
Figure 5: Adversarial robustness. ROC curves showing robust performance against three attack types with varying intensities.
Figure 6: Multi-bit scaling and information density. Watermark accuracy across different message bit lengths at fixed text length, demonstrating superior information density compared to multi-bit baselines with ¿90% accuracy up to 10 bits.
Figure 7: Framework component ablation studies. Left: Multi-bit watermarking accuracy scaling analysis with ablations. Right: ROC curves for feature concentration hyperparams ($\mu$, $\sigma$).
Understanding which implementation details drive SAEMark’s empirical success is critically important—our ablation studies validate that these practical optimizations are essential for bridging theoretical bounds and empirical performance, confirming our theoretical predictions about their necessity.
CheckAlignment filters.
The Range Similarity and Overlap Rate filters prove theoretically grounded and empirically validated. Figure 7 (left) demonstrates that the CheckAlignment algorithm’s 95% thresholds are not arbitrary—deviations cause significant degradation beyond 1,024 users (10 bits), confirming our theoretical analysis that these values optimally balance generation feasibility with discriminative power. These filters successfully compensate for theoretical independence assumptions when sequential generation creates dependencies in practice. Figure 7 (right) shows that the empirically-derived parameters achieve optimal ROC performance, validating our framework’s theoretical foundations.
Background feature masking.
This implementation detail proves essential for signal quality. Figure 12 in the appendix shows that removing the background mask causes AUC to plummet from  1.0 to 0.85. The mask excludes ubiquitous features like those related to punctuation or basic grammar patterns that would otherwise dominate FCS calculations without providing discriminative signals between different watermark keys. Detailed ablation results are provided in Appendix E.
These components work synergistically to enable SAEMark’s practical success: background masking isolates meaningful signals while alignment constraints makes watermark detection more accurate than the theoretical settings.
5 Conclusion
SAEMark introduces a fundamental paradigm shift in AI-generated content attribution through feature-based rejection sampling with sparse autoencoder representations. Our approach addresses critical limitations of existing watermarking methods by operating entirely through inference-time selection rather than model modification, enabling deployment with API-based services while maintaining superior text quality and detection accuracy. Three key advances enable this breakthrough: First, our general framework provides theoretical guarantees that relate watermark success probability to computational budget, independent of the specific feature extractor used. Second, the sparse autoencoder instantiation with Feature Concentration Scores captures meaningful semantic patterns that generalize across domains and languages. Third, practical optimizations including background feature masking and CheckAlignment filters bridge the gap between theoretical bounds and empirical performance, enabling efficient deployment. This work establishes that model interpretability tools can be effectively repurposed for content attribution tasks. The decoupling of watermarking from generation dynamics opens new possibilities for scalable, quality-preserving attribution systems that work seamlessly with existing language model APIs across diverse applications and languages.
Appendix A Limitations
We also found some limitations with our current approach. First, the method’s effectiveness depends on SAE feature quality. But be noted that this does not affect the applicability of our algorithm on the base LLMs, since we only apply SAEs on the Anchor LLM and require only access to the output texts from the base LLM, and we have a lot of pretrained SAEs from the open-source community that exhibit strong performance in interpreting model outputs. Second, detection watermarks effectively requires open-ended generation tasks, making attribution challenging for very short outputs like multiple-choice problems that only contain option keys. However this is a universal challenge for all watermarking algorithms, since short texts inevitably contains less information and less space to inject additional signatures.
These constraints reflect tradeoffs in privacy-preserving watermarking. Future work could explore dynamic candidate pruning to address these limitations. Nevertheless, our experiments across 4 benchmarks suggest these constraints pose manageable practical impacts compared to the system’s ethical advantages.
Appendix B Experimental Setup and Hyperparameter Details
This appendix provides a comprehensive description of the experimental setup, encompassing the hyperparameters and software configurations employed in this study.
B.1 Hyperparameters (SAEMark)
The following hyperparameters were used for the SAEMark:
  • Candidate Number (N): 50. This parameter denotes the number of candidate sequences sampled from the LLM.
  • Unit Number (M): 10. This specifies the number of discrete generation units produced by the model per attempt.
  • Attempt Number (K): 15. This metric represents the maximum times that the algorithm attempts to get an alignment.
B.2 Model Configuration
The section outlines the hyperparameter by the model during generation.
  • Base Model: Qwen2.5-7B-Instruct. This is the model on which the algorithm operates.
  • Sampling: This algorithm enables the model to generate various candidates, for which the parameter do_sampledo\_sample is set to $True$.
  • Temperature: This controls the randomness of the predictions by scaling the logits. The metric is set to $0.7$.
  • Max New Tokens: This specifies the maximum number of new tokens that the model can generate, which is $20$ during generation.
Appendix C Introduction to baselines
C.1 Single-bit Watermarks
KGW (Kirchenbauer et al., 2023)
The Key-based Green-list Watermarking (KGW) algorithm is a modern approach for watermarking text generated by LLMs. This method builds upon the work of (Kirchenbauer et al., 2023), who introduced a watermarking scheme that divides the token set into ’red’ and ’green’ lists based on a secret key and previously generated tokens.
Key features of KGW include the bifurcation of the token set into ’red’ and ’green’ lists, the use of a random seed dependent on a secret key and hash of prior tokens, reweighting of token log-probabilities to favor green tokens, and the introduction of permutation-based reweight strategies. These elements work in concert to create an effective watermarking system that balances detectability with output quality preservation.
The approach offers a balance between watermark embedding and preservation of text quality, addressing challenges faced by previous watermarking methods.
Unigram (Zhao et al., 2023)
The Unigram-Watermark and KGW algorithms, both designed for watermarking LLM-generated text, have distinct characteristics. Unigram-Watermark operates on individual tokens, using a consistent green list for each new token, while KGW employs a $K$-gram approach with varying green lists. Unigram-Watermark’s simplicity offers enhanced robustness against editing attacks and requires minimal implementation overhead. This streamlined approach leads to potential efficiency gains in both watermark embedding and detection processes, setting it apart from the more complex K-gram nature of KGW.
SWEET (Lee et al., 2024)
The Segment-Wise Entropy-based Embedding Technique (SWEET) is an innovative approach to watermarking code generated by large language models. SWEET addresses the challenge of maintaining code functionality while embedding detectable watermarks. It operates by selectively applying watermarking to high-entropy segments of the generated code, thereby preserving the overall code quality. This method significantly improves code quality preservation while outperforming baseline methods in detecting machine-generated code. SWEET achieves this by removing low-entropy segments during both the generation and detection of watermarks, effectively balancing the trade-off between detection capability and code quality degradation.
UPV (Liu et al., 2023)
The key feature of UPV is its use of separate neural networks for watermark generation and detection, addressing the limitation of shared key usage in previous methods. This separation allows for public verification without compromising the watermark’s security. UPV employs shared token embedding parameters between the generation and detection networks, enabling efficient and accurate watermark detection. The algorithm embeds small watermark signals into the LLM’s logits during generation, similar to existing methods, but uniquely conceals the watermarking details in the detection process. This approach ensures high detection accuracy while maintaining computational efficiency, and significantly increases the complexity of forging the watermark, thus enhancing its security in public detection scenarios.
DIP (Wu et al., 2023)
The Distribution-Preserving Watermarking (DIP) algorithm represents a significant advancement in watermarking techniques for large language models (LLMs). DIP’s innovation is its ability to maintain the original token distribution of the LLM while embedding a watermark, addressing a critical limitation of previous methods. This distribution-preserving property is achieved through a novel permutation-based approach that reweights token probabilities without altering the overall distribution. DIP offers provable guarantees on distribution preservation, detectability, and resilience against text modifications. The algorithm employs a texture key generation mechanism that considers multiple previous tokens, enhancing its robustness. Notably, DIP maintains text quality comparable to the original LLM output, owing to its distribution-preserving nature.
Unbiased (Hu et al., 2023b)
Unbiased watermarking and DIP watermarking are closely related concepts in the field of text watermarking for large language models (LLMs). Both approaches aim to embed watermarks while maintaining the original distribution of the LLM’s output. The key distinction lies in their theoretical foundations and implementation. Unbiased watermarking ensures that the expectation of the watermarked distribution matches the original distribution, while DIP watermarking guarantees that the watermarked distribution is identical to the original for every input. In essence, unbiased watermarking can be viewed as a relaxed version of DIP watermarking. While unbiased watermarking allows for small deviations in individual instances, DIP watermarking maintains strict distribution preservation. This relationship highlights a spectrum of watermarking techniques, where unbiased methods offer a balance between practicality and distribution preservation, while DIP methods provide stronger theoretical guarantees at potentially higher computational costs.
SynthID (Dathathri et al., 2024)
SynthID is an advanced watermarking method for large language models (LLMs) that builds upon previous work in generative text watermarking. The key innovation of SynthID lies in its use of Tournament sampling, which provides superior detectability compared to existing methods. This approach offers rigorous and customizable non-distortion properties, allowing for text quality preservation while maintaining effective watermarking. SynthID has been empirically validated, including through real user feedback from millions of chatbot interactions. Notably, the method introduces an algorithm to combine generative watermarking with speculative sampling, enabling efficient deployment in high-performance, large-scale production LLMs.
EXP (Aaronson & Kirchner, 2022)
EXP employs a pseudorandom function $f_{s}()$ with a secret seed $s$ known only to the model provider. Given previous tokens $w_{1},...,w_{t-1}$ and GPT’s probability distribution $p_{1},...,p_{K}$ for the next token $w_{t}$, the algorithm generates real values $r_{i}\in[0,1]$ using $f_{s}()$. EXP then selects the token $i$ that maximizes $r_{i}^{1/p_{i}}$. To detect the watermark, it calculates $\sum_{t=1}^{T}\ln\frac{1}{1-r_{t}^{\prime}}$ and compares it to a threshold. The scheme preserves the original token distribution while embedding a detectable watermark, with theoretical analysis showing distinct expected values for normal and watermarked text. The number of tokens required for reliable detection is $O(\frac{1}{\alpha^{2}}\log\frac{1}{\delta})$, where $\alpha$ is the average entropy per token and $\delta$ is the acceptable misclassification probability.
C.2 Multi-bit Watermarks
C.2.1 Baselines
CTWL (Wang et al., 2023a)
CTWL is a framework designed to embed multi-bit customizable information into texts produced by large language models (LLMs). It allows watermarks to carry details such as model version, generation time, and user ID. CTWL provides a mathematical model for watermarking and a comprehensive evaluation system that considers factors like success rate, robustness, coding rate, efficiency, and text quality. The Balance-Marking method uses a proxy language model to partition vocabulary probabilities, aiming to maintain watermarked text quality and achieve strong performance in evaluations. CTWL seeks to integrate multi-bit information watermarks into LLMs and offers a practical approach for tracing machine-generated texts.
Waterfall (Lau et al., 2024)
Waterfall is a training-free framework designed for robust and scalable text watermarking. It leverages large language models (LLMs) as paraphrasers to generate diverse text variations while preserving semantic meaning. By combining vocab permutation with orthogonal perturbation techniques, Waterfall aims to achieve scalability and robust verifiability while maintaining text fidelity. The framework supports multi-bit watermarks, enabling it to accommodate multiple users while ensuring effective watermark detection. Waterfall allows for a trade-off between watermark strength and text quality, making it adaptable to various requirements.
CODEIP (Yoo et al., 2023b)
CODEIP embeds a multi‑bit message into generated code by softly biasing token logits during decoding. At each step, a hash of the previous token and the secret ID selects “watermark” tokens whose logits $L_{WM}$ are boosted proportional to a strength $\beta$ . To guarantee syntactic validity, a pretrained type predictor assigns logits $L_{TP}$ only to tokens matching the expected lexical category, scaled by $\gamma$ . The final next‑token logits combine the base LLM scores, watermark bias, and grammar bias via
$w_{i}=\arg\max_{w\in V}\mathrm{softmax}\bigl{(}L_{LLM}+\beta L_{WM}+\gamma L_{TP}\bigr{)}$
Extraction recovers the ID by re‑hashing and finding the message maximizing cumulative watermark contributions.
REMARK-LLM (Zhang et al., 2024a)
REMARK-LLM introduces a robust watermarking framework for texts generated by large language models. It consists of three modules: message encoding, reparameterization, and message decoding. The message encoding module uses a sequence-to-sequence model to embed watermarks into LLM-generated texts. The reparameterization module applies Gumbel-Softmax to the dense token distribution into a sparser form. The message decoding module extracts watermarks using a transformer-based decoder. The framework incorporates malicious transformations during training to enhance robustness against attacks.
Provably Robust Multi-bit Watermarking for AI-generated Text (Qu et al., 2024)
The authors propose a multi‑bit watermarking scheme that embeds a user’s bit‑string ID into LLM‑generated text by first partitioning the message into pseudo‑randomly assigned bit‑segments per token, then biasing “green list” logits seeded by each segment’s value. A dynamic‑programming step balances token‑to‑segment assignments, and a Reed–Solomon error‑correction layer encodes segments to correct editing errors. Extraction enumerates only each segment’s possibilities $O(k\cdot 2^{b/k})$, yielding efficient, provably robust watermark recovery under bounded edit distance.
Robust Multi-bit Text Watermark with LLM-based Paraphrasers (Xu et al., 2024)
The authors propose a robust multi-bit text watermarking method using LLM-based paraphrasing. They design an encoder with two fine-tuned LLM paraphrasers ($\theta_{0}$ and $\theta_{1}$) that generate watermarked text by alternating based on the watermark code. A text segmentor divides the text into sentence-level segments, allowing each segment to encode one bit of the watermark message. The decoder uses a trained text classifier to determine the watermark bit for each segment. The method employs a co-training framework where the encoder and decoder are alternately updated. The decoder acts as a reward model during PPO-based reinforcement learning to fine-tune the encoder, optimizing both the detection of the watermark and the semantic similarity to the original text.
Robust Multi-bit Natural Language Watermarking through Invariant Features (Yoo et al., 2023a)
The authors propose a robust multi-bit natural language watermarking method based on invariant features. They identify key words and syntactic dependencies as invariant features to embed watermarks, leveraging these features’ resistance to minor textual modifications. A corruption-resistant infill model is also introduced to enhance watermark extraction robustness. Their method first selects mask positions based on these invariant features and then generates watermarked texts using an infill model. A robust infill model is developed to improve recovery of watermarked texts from corrupted versions.
C.2.2 Why Choose Waterfall as the Multi-Bit Baseline?
Among the various text watermarking methods, Waterfall offers distinct advantages for multi-bit applications. As a training-free framework, it efficiently generates diverse text variations using LLMs as paraphrasers while preserving semantic meaning. Its key advantages include complexity that doesn’t depend on word or sentence count, allowing scalability. It also offers evaluation metrics akin to single-bit methods and supports multi-language and multi-dataset watermarking, making it highly adaptable.
Figure 8: An example of Feature Concentration Score (FCS) calculation process.
Appendix D Details of FCS Generation
This section elaborates on the methodology behind the generation of the Feature Concentration Score (FCS). The process is illustrated in Figure 8, which outlines four key steps.
Input: Token sequence $T$, SAE $\theta$ for the entire sequence
Output: Feature Concentration Score (FCS)
$\Phi\leftarrow\theta(T)$, yielding activation vectors $\phi_{1},\phi_{2},...,\phi_{n}$ for each token position in $T$;
$indices\leftarrow[]$;
for $t=1$ to $n$ do
 $\phi_{t}$ is the activation vector for token at position $t$;
 $index\leftarrow\arg\max_{i}(\phi_{t}\odot m)_{i}$;
   Append indexindex to indicesindices;
end for
$featureSet\leftarrow set(indices)$, removing duplicates;
$featureSum\leftarrow 0$;
$totalNorm\leftarrow 0$;
for $t=1$ to $n$ do
 $tokenSignificance\leftarrow 0$;
 foreach $i\in featureSet$ do
    $tokenSignificance\leftarrow tokenSignificance+\phi_{t,i}$;
   end foreach
 $featureSum\leftarrow featureSum+tokenSignificance$;
[…]
 // Accumulate significant features and norms
end for
$FCS\leftarrow\frac{featureSum}{totalNorm}$;
// Calculate final FCS
return FCSFCS;
[…]
Extracting SAE Features for Each Token
Given a token sequence $T$, we utilize SAE to derive an activation vector $\phi_{t}$ for each token position $t$. This vector, $\phi_{t}$, embodies the representation of the token at position $t$ with a dimensionality of 16,384.
Selecting the Most Significant Feature
For every activation vector $\phi_{t}$, our objective is to identify the most significant feature, which serves as a descriptor for the token at position $t$. This is achieved through applying the function $argmax_{i}(\phi_{t}\odot m)_{i}$, where $m$ is a mask. The output of this function yields the indices corresponding to the most prominent feature, denoted as ”SAE Feature Indices” in Figure  Figure 8.
Aggregating Most Significant Features
As depicted in Figure  Figure 8, each token’s position $t$ has its most significant feature. However, when summarizing the critical features of the entire sequence $T$, redundancies may occur. To address this, we employ a set operation to eliminate duplicate entries among the significant features, resulting in a unique collection termed as ”Feature Set SS”.
Calculating Feature Concentration Score
Upon obtaining the Feature Set $S$, we aim to quantify how these significant features contribute to the overall sequence $T$ concerning SAE feature values. For each $\phi_{t}$, we compute the sum of $\phi_{t,i}$, where $i$ represents the index belonging to $S$. This aggregate score measures the contribution of significant features to individual tokens within $T$. Accumulating this metric across all tokens provides a global measure for the sequence.
To evaluate the total activation value of SAE features over the sequence $T$, we apply the $L1$ norm to each $\phi_{t}$, obtaining the sum of absolute values for each token’s feature vector. Summing these across all tokens yields the total SAE value for $T$. The Feature Concentration Score (FCS) is defined as the ratio of the accumulated contributions of significant features to the total SAE feature values.
The detailed steps for computing the FCS are outlined in algorithm 3.
This score effectively captures the concentration of key features within a token sequence and is useful for applications in watermark embedding.
Appendix E Additional Experimental Results
E.1 Adversarial Robustness Evaluation
Word Deletion Attack
In the main text, we conducted experiments using the ”maintain content structure” version of the word deletion attack. However, the original word deletion attack involves splitting a paragraph and randomly removing words, which disrupts the structure that watermarking methods rely on, making it harder for the detection system to identify the watermark. To address this issue, we modified the attack to preserve structure while still performing word deletions. By maintaining the integrity of the structure, the attack bypasses watermark detection more effectively.
In our experimental results, we compare two versions of the word deletion attack. The ”keep structure” method, represented in a darker color, shows more robust performance with higher AUC values (0.949 at $\epsilon=0.05$ and 0.858 at $\epsilon=0.1$). In contrast, the ”not keep structure” method, shown in a lighter color, demonstrates a decline in performance, with AUC values dropping to 0.901 at $\epsilon=0.05$ and 0.825 at $\epsilon=0.1$. These results indicate that preserving the content structure during the attack strengthens the watermark’s resistance, whereas random word deletions that disrupt the structure reduce detection accuracy.
As shown in the Figure 9, the ”keep structure” method outperforms the ”not keep structure” method in terms of AUC, demonstrating its effectiveness in watermark resistance.
Figure 9: Word Deletion on SAEMark ROC curves highlighting the performance difference between ”keep structure” and ”not keep structure” methods under word deletion attacks with varying intensities (5%, 10%).
Basic Synonym Substitution Attack
Our study also examines ”keeping structure” versus ”not keeping structure” approaches in the context of basic synonym substitution attacks, which are less likely to disrupt the content’s structural integrity.
Figure 10: Basic Synonym Substitution on SAEMark ROC curves comparing ”keep structure” and ”not keep structure” methods under basic synonym substitution attacks at different intensities (5%, 10%).
Figure 10 shows ROC curves comparing model performance under different conditions, with the original non-structure-preserving method in lighter shades and the modified structure-preserving method in darker hues. The analysis reveals minimal differences in AUC values between the two, indicating similar model resilience to both forms of synonym substitution. Notably, the model demonstrates performance robustness that exceeds that observed in deletion attack scenarios, reflected by AUC scores that remain close to the baseline.
Context-aware Synonym Substitution Attack
Due to our algorithm’s prominent performance against context-aware synonym attack. More intensities (20%, 30%, 40%, 50%) are carried upon these kinds of attacking.
Figure 11: Context-Aware Synonym Substitution on SAEMark ROC curves comparing ”keep structure” and ”not keep structure” methods under basic synonym substitution attacks at different intensities (5%, 10%).
The results of the context-aware watermarking method, shown in  Figure 11 tested under this attack, demonstrate substantial robustness. Even with high substitution ratios—up to 50% token replacement—the AUC remains relatively high, highlighting the method’s ability to maintain detection performance under significant adversarial pressure. The ROC curves further corroborate this, showing that the true positive rate remains consistently high across varying false positive levels, even as attack intensity increases. This demonstrates a well-balanced trade-off between true and false positives, ensuring reliable detection without excessive false alarms. These findings affirm that the watermarking method is both effective and robust, offering reliable protection against sophisticated attacks while maintaining strong detection accuracy.
E.2 Ablation Study on Background Frequent Features
In section 3, we utilize $\phi_{t}\odot m$, where $m$ is a mask that excludes background frequent features.
In this section, we generate the Feature Concentration Score (FCS) without using $m$ and conduct ROC experiments for further analysis. To evaluate the impact of background frequent feature masking on our model’s performance, we performed an ablation study.
With background frequent feature masking in place, the model achieved an AUC of almost 1.0. Upon removing this masking, the AUC dropped to 0.85, as illustrated in Figure 12. This significant decrease demonstrates that background frequent feature masking plays a crucial role in our algorithm, emphasizing its importance for optimal performance.
Figure 12: Ablation on Background Frequent Feature Masking The ROC curve compares the performance with and without background frequent feature masking.
Appendix F Use Of AI Assistants
We employed AI assistants for two tasks: (1) generating routine code implementations and boilerplate functions, and (2) performing grammatical review and sentence-level editing of the manuscript. All AI-generated content underwent thorough manual review. The core research methodology, findings, and analysis remain entirely our own work.

### v2

@@
-SAEMark: Multi-bit LLM Watermarking with Inference-Time Scaling
+SAEMark: Steering Personalized Multilingual LLM Watermarks with Sparse Autoencoders
 Abstract
-Watermarking LLM-generated text is critical for content attribution and misinformation prevention.
-However, existing methods compromise text quality, require white-box model access and logit manipulation—limitations that exclude API-based models and multilingual scenarios.
-We propose SAEMark, a general framework for post-hoc multi-bit watermarking that embeds personalized messages solely via inference-time, feature-based rejection sampling without altering model logits or requiring training.
-Our approach operates on deterministic features extracted from generated text, selecting outputs whose feature statistics align with key-derived targets.
-This framework naturally generalizes across languages and domains while preserving text quality through sampling LLM outputs instead of modifying.
-We provide theoretical guarantees relating watermark success probability and compute budget that hold for any suitable feature extractor; empirically, we demonstrate the framework’s effectiveness using Sparse Autoencoders (SAEs), achieving superior detection accuracy and text quality.
-Experiments across 4 datasets show SAEMark’s consistent performance, with 99.7% F1 on English and strong multi-bit detection accuracy.
-SAEMark establishes a new paradigm for scalable watermarking that works out-of-the-box with closed-source LLMs while enabling content attribution.
-[^1] We open-source code and data at: https://zhuohaoyu.github.io/SAEMark
+Watermarking LLM-generated text is critical for content attribution and misinformation prevention, yet existing methods compromise text quality and require white-box model access with logit manipulation or training, which exclude API-based models and multilingual scenarios.
+We propose SAEMark, an inference-time framework for multi-bit watermarking that embeds personalized information through feature-based rejection sampling, fundamentally different from logit-based or rewriting-based approaches: we do not modify model outputs directly and require only black-box access, while naturally supporting multi-bit message embedding and generalizing across diverse languages and domains.
+We instantiate the framework using Sparse Autoencoders as deterministic feature extractors and provide theoretical worst-case analysis relating watermark accuracy to computational budget.
+Experiments across 4 datasets demonstrate strong watermarking performance on English, Chinese, and code while preserving text quality.
+SAEMark establishes a new paradigm for scalable, quality-preserving watermarks that work seamlessly with closed-source LLMs across languages and domains.
 1 Introduction
-Large language models (LLMs) have revolutionized text generation across domains, from creative writing to code synthesis (Brown et al., 2020; Guo et al., 2024).
+Figure 1: An overview of SAEMark.
+Large language models (LLMs) have revolutionized text generation across domains, from creative writing to code synthesis [6, 10].
 However, their ability to produce human-quality text at scale raises serious concerns about misinformation, copyright infringement, and content laundering.
 As these models become ubiquitous, reliably attributing AI-generated content becomes critical for accountability and trust.
-Watermarking—embedding detectable signatures into generated text—offers a promising solution, but existing approaches face a fundamental tradeoff.
+Watermarking—embedding detectable signatures into generated text—offers a promising solution.
 They must preserve text quality while enabling reliable detection, operate across languages and domains, and scale to distinguish between many users or sources.
@@
 Existing watermarking methods struggle with these requirements.
-Token-level approaches like KGW (Kirchenbauer et al., 2023) and EXP (Aaronson & Kirchner, 2022) require direct access to model logits, excluding API-based deployment, and can degrade text quality through probability manipulation.
-Syntactic methods (Hou et al., 2023) fail to generalize across languages, while specialized approaches (Lee et al., 2024) work well in narrow domains but break down when applied more broadly.
-Even recent black-box methods (Bahri & Wieting, 2024; Chang et al., 2024) rely on surface-level statistics or require auxiliary models, limiting their robustness and scalability.
+Token-level approaches like KGW [16] and EXP [1] require direct access to model logits, excluding API-based deployment, and can degrade text quality through probability manipulation.
+Syntactic methods [11] fail to generalize across languages, while specialized approaches [18] work well in narrow domains but break down when applied more broadly.
+Even recent black-box methods [4, 7] rely on surface-level statistics or require auxiliary models, limiting their robustness and scalability.
 We introduce SAEMark, a fundamentally different approach that sidesteps these limitations entirely.
@@
 Our contributions span theory and practice.
-We develop a general framework for watermarking through feature-guided selection that works with any feature extractor and any language model API.
-We provide theoretical guarantees that explain how watermark success scales with computational resources and text length, independent of the specific features used.
-Finally, we demonstrate a practical instantiation using Sparse Autoencoders that achieves superior detection accuracy and text quality across English, Chinese, and code, encoding more information per unit length than existing multi-bit approaches.
-Figure 1: An overview of SAEMark.
-2 Preliminaries
-2.1 Related Work
-LLM watermarking is a technique to embed special patterns into the output of LLMs, and has traditionally been used to identify LLM generated text from human-written text (Jawahar et al., 2020).
-Different from post-hoc detection methods (Zellers et al., 2019) that analyze statistical patterns in existing text, language model watermarking aims to embed detectable signatures during generation (Kirchenbauer et al., 2023).
-Existing approaches exhibit several limitations.
-Many methods compromise generation quality through direct manipulation of token probabilities (Kirchenbauer et al., 2023) or syntactic modifications (Atallah et al., 2002).
-The challenge of language and domain generalization remains largely unaddressed, with current techniques primarily optimized for English and struggling with multilingual content or specialized domains like code (Lee et al., 2024).
-Notably, PersonaMark (Zhang et al., 2024b) represents early attempts at personalized watermarking, but its reliance on English-specific syntactic patterns and closed-source implementation makes scalability and cross-lingual capability difficult to verify.
-Recently, more multi-bit watermarking methods have been proposed to embed multiple bits of information into generated text (Lau et al., 2024; Zhang et al., 2024a; Yoo et al., 2023a; Guan et al., 2024; Yoo et al., 2023b; Qu et al., 2024), primarily by extending single-bit watermarking that manipulates logits during generation; these methods inherit the limitations of single-bit designs.
-Complementary black-box watermarks avoid white-box logit access by using post-hoc selection or rewriting (Chang et al., 2024).
-However, they typically operate on surface statistics or introduce auxiliary model dependencies and do not directly address multi-bit message embedding at scale.
-Our framework differs by performing inference-time selection among naturally generated candidates using deterministic feature statistics computed over domain-appropriate units.
+We develop a general framework for watermarking through feature-guided selection that works with any feature extractor and LLM.
+We provide theoretical guarantees that explain how SAEMark accuracy scales with compute budget, independent of feature extractors.
+Finally, we demonstrate a practical instantiation using SAE as feature extractor that achieves superior accuracy and text quality across languages and domains, encoding more information than existing approaches.
+2 Related Work
+LLM watermarking is a technique to embed special patterns into the output of LLMs, and has traditionally been used to identify LLM generated text from human-written text [14].
+Different from post-hoc detection methods [31] that analyze statistical patterns in existing text, language model watermarking aims to embed detectable signatures during generation [16].
+These methods compromise generation quality through direct manipulation of token probabilities [16] or syntactic modifications [2].
+The challenge of language and domain generalization remains largely unaddressed, with current techniques primarily optimized for English and struggling with multilingual content or specialized domains like code [18].
+Notably, PersonaMark [33] represents early attempts at personalized watermarking, but its reliance on English-specific syntactic patterns and closed-source implementation makes scalability and cross-lingual capability difficult to verify.
+Recently, more multi-bit watermarking methods have been proposed to embed multiple bits of information into generated text [17, 32, 29, 9, 30, 22], primarily by extending single-bit watermarking that manipulates logits during generation; these methods inherit the limitations of single-bit designs.
+Complementary black-box watermarks avoid white-box logit access by using post-hoc selection or rewriting [7].
+However, they typically operate on statistics or introduce auxiliary model dependencies and do not directly address multi-bit message embedding at scale.
+Our framework differs by performing inference-time selection among naturally generated candidates using deterministic feature statistics.
 This enables extractor-agnostic analysis and multilingual, domain-agnostic multi-bit watermarking without modifying model logits.
-2.2 Sparse Autoencoders
-Sparse Autoencoders (SAEs) are pre-trained interpretability tools that decompose LLM activations into human-understandable features (Bricken et al., 2023).
+Sparse Autoencoders (SAEs) are pre-trained interpretability tools that decompose LLM activations into human-understandable features [5].
 For a given base model $\mathcal{M}$ and layer $l$, an SAE processes hidden states $\mathbf{h}_{t}$ at position $t$ as:
@@
 $\mathcal{L}=\underbrace{\|\mathbf{h}_{t}-\text{Dec}(\mathbf{f}_{t})\|^{2}}_{L_{rec}}+\lambda\underbrace{\|\mathbf{f}_{t}\|_{1}}_{L_{sparse}}$ (2)
-This training produces features that correspond to interpretable concepts (Bricken et al., 2023; Templeton et al., 2024).
-For instance, SAEs applied to Gemma 2B (Team et al., 2024)’s final transformer layer, GemmaScope, reveal features like ”Python function definitions” or ”Concept related to color blue” (Lieberum et al., 2024).
-Our watermarking leverages three key properties of pre-trained SAE features.
-First, layer-specific patterns capture distinct behaviors from different model layers.
-Second, multilingual activation allows the same features to fire for equivalent concepts across languages.
-Third, sparsity enables efficient analysis through few active features per token.
+This training produces features that correspond to interpretable concepts [5, 24] like “Function definitions” or “Concept related to color blue” [19].
+Our watermarking leverages key properties of pre-trained SAEs: multilingual activation allows the same features to fire for equivalent concepts across languages deterministically.
+sparsity enables efficient analysis through few active features per token.
 These properties support language-agnostic statistics via masked feature aggregations:
 $\phi(\mathbf{y})=\frac{1}{|\mathbf{y}|}\sum_{t}\mathbf{f}_{t}\odot\mathbf{m}$ (3)
-where $\mathbf{m}$ filters background features that fire ubiquitously regardless of content (e.g., punctuation-associated features).
-The summary $\phi(\mathbf{y})$ provides a deterministic statistic used by our generation and detection procedures (Sec 3).
+where $\mathbf{m}$ filters background features that fire ubiquitously regardless of content (e.g., punctuation).
+The summary $\phi(\mathbf{y})$ provides a deterministic statistic used by our watermarking procedures (Sec 3).
+3 Methodology
+We present a general framework for post-hoc, multi-bit watermarking via feature-based rejection sampling.
+The key observation is that different LLM generations produce distinct values of deterministic feature statistics computed over domain-appropriate units, and these statistics can be steered by selecting among naturally generated candidates, without modifying model logits, parameters or generated texts.
+We structure the section as follows: (1) a general framework that is extractor-agnostic, (2) theoretical guarantees with an emphasis on worst-case bounds, and (3) an effective instantiation using sparse autoencoders as feature extractors.
 3.1 Task Definition
@@
 Given a base LLM $\mathcal{M}$, watermark key $k\in\mathcal{K}$, input $\mathbf{x}$, and message $m\in\{0,1\}^{b}$, the algorithm produces $\mathbf{y}$ by post-hoc selection over $\mathcal{M}$’s outputs (black-box/API-compatible; no access to logits or parameters; cf.
-black-box watermarking (Bahri & Wieting, 2024; Chang et al., 2024)):
+black-box watermarking [4, 7]):
 $\mathbf{y}=\text{Mark}(\mathcal{M},\mathbf{x},k,m).$ (4)
@@
 […]
-The scheme targets three properties: key privacy (deriving $k$ from watermarked outputs is hard), verifier-held detectability (any party holding $k$ can verify), and collusion resistance (multiple keys should not facilitate removal or forgery).
-Threat model and scope.
-Our focus is attribution without storing sensitive content.
-This work does not claim cryptographic unforgeability when keys are known; preventing adversarial forgeries under black-box constraints is an important direction for security-focused follow-ups.
-3 Methodology
-We present a general framework for post-hoc, multi-bit watermarking via feature-based rejection sampling.
-The key observation is that different LLM generations produce distinct values of deterministic feature statistics computed over domain-appropriate units, and these statistics can be steered by selecting among naturally generated candidates, without modifying model logits, parameters or generated texts.
-We structure the section as follows: (1) a general framework that is extractor-agnostic, (2) theoretical guarantees with an emphasis on worst-case bounds, and (3) an effective instantiation using sparse autoencoders as feature extractors.
+Threat model and scope: The scheme targets three properties: key privacy (deriving $k$ from watermarked outputs is hard), verifier-held detectability (any party holding $k$ can verify), and collusion resistance (multiple keys should not facilitate forgery).
+Our focus is attribution without storing LLM generated text.
+This work does not claim cryptographic unforgeability when keys are known; preventing adversarial forgeries is an important direction for security-focused follow-ups.
 3.2 General Framework for Feature-based Watermarking
@@
 This process steers generation toward key-dependent patterns without modifying the underlying language model.
+Input: Prompt $c$, key $k$, LLM $G$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, units $M$, attempts $K$, candidates NN
+Output: Watermarked text $x^{*}$
+for $attempt\leftarrow 1$ to $K$ do
+  $x^{*}\leftarrow c$; $\{\tau_{i}\}_{i=1}^{M}\leftarrow\text{TargetsFromKey}(k)$; $\{z_{i}\}\leftarrow\emptyset$
+  for $i\leftarrow 1$ to $M$ do
+     $\mathcal{X}\leftarrow\text{GenerateCandidates}(G,x^{*},N)$
+     $x_{best}\leftarrow\arg\min\limits_{x\in\mathcal{X}}\big|\hat{F}(s(\phi(x))){-}\tau_{i}\big|$
+     $x^{*}\leftarrow x^{*}\oplus x_{best}$; $z_{i}\leftarrow\hat{F}(s(\phi(x_{best})))$
+  end for
+ if CheckAlignment($\{\tau_{i}\},\{z_{i}\}$) then
+    return $x^{*}$
+  end if
+end for
+return $x^{*}$
+Algorithm 1 Watermark generation
+Input: Text $x$, candidate keys $\mathcal{K}$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, alignment thresholds, significance $\alpha$
+Output: Detection result $d\in\mathcal{K}\cup\{\emptyset\}$
+$\{z_{j}\}\leftarrow[\hat{F}(s(\phi(u)))\;\forall u\in\text{SegmentByDomain}(x)]$
+$\mathcal{D}\leftarrow\emptyset$
+foreach $k_{i}\in\mathcal{K}$ do
+  $\{\tau_{j}\}\leftarrow\text{TargetsFromKey}(k_{i},|\{z_{j}\}|)$
+  if CheckAlignment($\{\tau_{j}\},\{z_{j}\}$) then
+     $t,p\leftarrow\text{StudentTTest}(\{z_{j}\},\{\tau_{j}\})$
+     if $t>t_{\alpha/2}\land p<\alpha$ then
+        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(k_{i},t)\}$
+     end if
+  end if
+end foreach
+return $\operatorname{argmax}_{(k_{i},t_{i})\in\mathcal{D}}t_{i}$ if $\mathcal{D}\neq\emptyset$ else $\emptyset$
+Algorithm 2 Watermark detection
+Figure 2: Pseudocode for SAEMark: generation and detection.
 Text segmentation.
@@
 $R_{\min}<\frac{z_{\max}-z_{\min}}{\tau_{\max}-\tau_{\min}}<R_{\max}$ (6)
-where $z_{\max}=\max_{i}z_{i}$, $z_{\min}=\min_{i}z_{i}$, and similarly for $\tau$. We typically set $R_{\min}=0.95$, $R_{\max}=1.05$.
+where $z_{\max}=\max_{i}z_{i}$, $z_{\min}=\min_{i}z_{i}$, and similarly for $\tau$. We set $R_{\min}=0.95$, $R_{\max}=1.05$.
 Overlap Rate Filter: This constraint ensures sufficient overlap between the value ranges of both sequences:
@@
 The key with the highest significance score is returned if it passes the threshold; otherwise, we classify the text as unwatermarked.
-Input: Prompt $c$, key $k$, LLM $G$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, units $M$, attempts $K$, candidates NN, alignment thresholds
-Output: Watermarked text $x^{*}$
-for $attempt\leftarrow 1$ to $K$ do
- $x^{*}\leftarrow c$; $\{\tau_{i}\}_{i=1}^{M}\leftarrow\text{TargetsFromKey}(k)$; $\{z_{i}\}\leftarrow\emptyset$
- for $i\leftarrow 1$ to $M$ do
-    $\mathcal{X}\leftarrow\text{GenerateCandidates}(G,x^{*},N)$
-    $x_{best}\leftarrow\arg\min\limits_{x\in\mathcal{X}}\big{|}\hat{F}(s(\phi(x))){-}\tau_{i}\big{|}$
-    $x^{*}\leftarrow x^{*}\oplus x_{best}$; $z_{i}\leftarrow\hat{F}(s(\phi(x_{best})))$
-   end for
- if CheckAlignment($\{\tau_{i}\},\{z_{i}\}$then
-    return $x^{*}$
-   end if
-end for
-return $x^{*}$
-Algorithm 1 Watermark generation
-Input: Text $x$, candidate keys $\mathcal{K}$, extractor $\phi$, statistic $s(\cdot)$ with CDF estimate $\hat{F}$, alignment thresholds, significance level $\alpha$
-Output: Detection result $d\in\mathcal{K}\cup\{\emptyset\}$
-$\{z_{j}\}\leftarrow[\hat{F}(s(\phi(u)))\;\forall u\in\text{SegmentByDomain}(x)]$
-$\mathcal{D}\leftarrow\emptyset$
-foreach $k_{i}\in\mathcal{K}$ do
- $\{\tau_{j}\}\leftarrow\text{TargetsFromKey}(k_{i},|\{z_{j}\}|)$
- if CheckAlignment($\{\tau_{j}\},\{z_{j}\}$then
-    $t,p\leftarrow\text{StudentTTest}(\{z_{j}\},\{\tau_{j}\})$
-    if $t>t_{\alpha/2}\land p<\alpha$ then
-       $\mathcal{D}\leftarrow\mathcal{D}\cup\{(k_{i},t)\}$
-      end if
-   end if
-end foreach
-return $\operatorname{argmax}_{(k_{i},t_{i})\in\mathcal{D}}t_{i}$ if $\mathcal{D}\neq\emptyset$ else $\emptyset$
-Algorithm 2 Watermark detection
-Figure 2: Pseudocode for SAEMark: generation and detection.
 3.3 Theoretical Analysis and Guarantees
@@
 $\mathbb{P}(\exists j:|S_{j}-\tau|\leq k\tau)\geq 1-(1-p_{\min})^{N}$ (8)
-Worst-case analysis and tight bounds.
+Worst-case analysis and bounds.
 To derive conservative guarantees, consider the worst-case target $\tau=\mu+2\sigma$ at the boundary of the feasible range. The single-candidate success probability becomes:
 $p_{\min}=\mathbb{P}((1-k)\tau\leq S_{j}\leq(1+k)\tau)=\Phi\left(\frac{(1+k)(\mu+2\sigma)-\mu}{\sigma}\right)-\Phi\left(\frac{(1-k)(\mu+2\sigma)-\mu}{\sigma}\right)$ (9)
-where $\Phi$ denotes the standard normal CDF. This simplifies to $p_{\min}=\Phi(2(1+k)+k\mu/\sigma)-\Phi(2(1-k)-k\mu/\sigma)$.
+where $\Phi$ denotes the standard normal CDF and $p_{\min}=\Phi(2(1+k)+k\mu/\sigma)-\Phi(2(1-k)-k\mu/\sigma)$.
 The fundamental insight is that observed feature statistics are tightly bounded to target values.
@@
 Sparse autoencoders—interpretability tools designed to understand language model internals—provide an ideal solution.
-They decompose language representations into interpretable semantic components (”technical writing,” ”narrative voice,” ”mathematical reasoning”) that exhibit distinctly different activation patterns across generations.
-By applying the sparse autoencoder to a separate ”anchor” model, our approach remains compatible with any target language model, including API services, while extracting the discriminative yet predictable statistics our framework requires.
+They decompose language representations into interpretable semantic components (e.g.
+“technical writing,” “math symbols”) that exhibit distinctly different activation patterns across generations.
+By applying the SAE to a separate “anchor” model, our approach remains compatible with any target language model, including API services, while extracting the discriminative yet predictable statistics our framework requires.
 Figure 3: Distribution analysis of FCS. FCS distribution with density estimation (left) and Q-Q plot (right); statistical tests support approximate normality.
@@
 Second, background feature masking ensures FCS calculations focus on discriminative semantic patterns rather than ubiquitous surface features.
-We precompute a mask excluding ”background” SAE features, like those related to punctuation or basic grammar, to focus on discriminative semantic patterns.
+We precompute a mask excluding "background" SAE features, like those related to punctuation or basic grammar, to focus on discriminative semantic patterns.
 With empirically observed parameters $\mu=0.142$, $\sigma=0.029$, and tolerance $\epsilon=0.1$, our bounds yield concrete success probabilities: $N=50$ achieves $99\%$ per-unit success, $N=20$ maintains $85.32\%$, and $N=10$ achieves $61\%$.
@@
 Table 1: Dataset Statistics. Characteristics of the multilingual benchmarks used in evaluation.
-C4 (2020) LCSTS (2015) MBPP (2021) PandaLM (2023b)
+C4 [23] LCSTS [12] MBPP [3] PandaLM [26]
 # Samples 500 500 257 169
 Language English Chinese Python English
-Task Type
-Completion
-Summarization
-Code Gen.
-Inst. Following
+Task Type Completion Summarization Code Generation Instruction Following
 From test split of sanitized version of MBPP.
-Setup.
 We evaluate on 4 diverse datasets as shown in Table 1.
@@
 For text quality, we report win-rates of pairwise comparison on PandaLM judged by GPT-4o in our main results, and average pointwise scores on BIGGen-Bench judged by their officially released judge model as an alternative text-quality experiment.
-We use implementation for single-bit watermarks from MarkLLM (Pan et al., 2024) toolkit and Waterfall (Lau et al., 2024) as it’s the current best open-source training-free multi-bit watermark similar to our setting.
+We use implementation for single-bit watermarks from MarkLLM [21] toolkit and Waterfall [17] as it’s the current best open-source training-free multi-bit watermark similar to our setting.
 Full details of baselines in Appendix C.
@@
 All metrics are reported as percentages (%).
-Method C4 (English, 2020) LCSTS (Chinese, 2015) MBPP (Code, 2021) PandaLM (Instruction,  2023b)
+Method C4 (English, [23])     LCSTS (Chinese, [12]) MBPP (Code, [3])     PandaLM (Instruction,  [26])
 […]
 Single-bit Watermarks
-KGW (2023) 99.2 99.6 99.2 99.1 98.8 99.1 65.4 31.9 48.0 41.5 89.9 80.4 88.8
-EXP (2022) 99.5 99.6 99.5 99.3 99.4 99.3 57.8 16.7 28.4 23.2 79.3 59.4 74.2
-UPV (2023) 86.0 72.0 83.7 90.5 91.0 90.5 51.6 3.1 6.0 36.0 54.0 8.0 14.8
-Unigram (2023) 98.8 98.6 98.8 98.2 97.0 98.2 65.4 31.9 48.0 35.3 53.3 7.2 13.4
-DIP (2023) 96.0 92.6 95.9 97.7 96.2 97.7 60.7 22.6 36.5 36.5 81.5 63.8 77.5
-Unbiased (2023b) 96.7 94.4 96.6 97.8 96.4 97.8 64.0 29.2 44.8 40.2 74.3 49.3 65.7
-SynthID (2024) 98.2 97.2 98.2 97.6 96.2 97.6 62.5 26.1 41.0 36.0 81.2 63.0 77.0
-SWEET (2024) 99.6 99.6 99.6 50.0 0.0 0.0 72.4 45.9 62.4 47.2 87.7 76.8 86.2
+KGW [16] 99.2 99.6 99.2 99.1 98.8 99.1 65.4 31.9 48.0 41.5 89.9 80.4 88.8
+EXP [1] 99.5 99.6 99.5 99.3 99.4 99.3 57.8 16.7 28.4 23.2 79.3 59.4 74.2
+UPV [20] 86.0 72.0 83.7 90.5 91.0 90.5 51.6 3.1 6.0 36.0 54.0 8.0 14.8
+Unigram [34] 98.8 98.6 98.8 98.2 97.0 98.2 65.4 31.9 48.0 35.3 53.3 7.2 13.4
+DIP [27] 96.0 92.6 95.9 97.7 96.2 97.7 60.7 22.6 36.5 36.5 81.5 63.8 77.5
+Unbiased [13] 96.7 94.4 96.6 97.8 96.4 97.8 64.0 29.2 44.8 40.2 74.3 49.3 65.7
+SynthID [8] 98.2 97.2 98.2 97.6 96.2 97.6 62.5 26.1 41.0 36.0 81.2 63.0 77.0
+SWEET [18] 99.6 99.6 99.6 50.0 0.0 0.0 72.4 45.9 62.4 47.2 87.7 76.8 86.2
 Multi-bit Watermarks
-Waterfall (2024) 93.6 88.0 93.2 95.3 91.6 95.1 52.5 6.2 11.6 46.4 73.2 47.1 63.7
+Waterfall [17] 93.6 88.0 93.2 95.3 91.6 95.1 52.5 6.2 11.6 46.4 73.2 47.1 63.7
 SAEMark (Ours) 99.7 99.8 99.7 99.2 99.6 99.2 74.5 50.2 66.3 67.6 86.6 73.9 84.6
@@
 Beyond accuracy, Table 2 shows SAEMark achieves the highest quality score (67.6%) on PandaLM as judged by GPT-4o pairwise comparisons.
-To study how this conclusion generalizes across different backbone models, since backbone performance is the most significant factor affecting text quality, we conduct additional evaluation on BIGGen-Bench comparing against both watermarked baselines and unwatermarked text.
-Table 3: Text quality evaluation on BIGGen-Bench. Scores are on a 5-point Likert scale (higher is better), judged by the officially released BIGGen-Bench judge model (Kim et al., 2024).
+To study how this generalizes across different backbone LLMs, we conduct additional evaluation on BIGGen-Bench comparing against watermarked baselines and unwatermarked text.
+Table 3: Text quality evaluation on BIGGen-Bench. Scores are on a 5-point Likert scale (higher is better) [15].
 Model Unwatermarked SAEMark KGW Waterfall
@@
 gemma-3-4b-it 4.26 4.23 3.98 4.19
-Table 3 confirms SAEMark consistently achieves the highest quality among watermarking methods across three backbone models.
-This quality advantage stems from a fundamental difference in approach: rather than manipulating logits or applying external rewriting to obtain watermarked text, we simply select among naturally generated candidates.
-This post-hoc selection guarantees that every watermarked segment remains an unmodified LLM generation from the backbone model itself, ensuring text quality stays bounded by the its own capabilities.
+Table 3 confirms SAEMark achieves the highest quality among watermarks across three backbone LLMs.
+This quality advantage stems from our design: rather than manipulating logits or applying external rewriting to obtain watermarked text, we simply run post-hoc selection among naturally generated candidates, ensuring text quality stays bounded by the its own capabilities.
 4.3 Computational Overhead and Scalability
@@
 N=5 N=10 N=20 N=50
-C4 (Raffel et al., 2020)
+C4 [23]
 Acc. 98.7 99.2 98.7 99.7
@@
 F1 86.8 98.0 98.7 99.7
-LCSTS (Hu et al., 2015)
+LCSTS [12]
 Acc. 98.6 99.0 98.6 99.2
@@
 F1 83.6 97.5 98.6 99.2
-(b) Perf. vs. Average Latency
+(b) Perf. vs. End to end Latency
 Method Acc. Rec. F1 Latency
@@
 In contrast, these optimized frameworks do not provide efficient watermark implementations for logit-manipulation methods, as such implementations require significant backend rewriting and may impact performance.
-While this creates a significant latency difference despite some methods theoretically needing less compute overhead, we consider this a practical advantage reflecting the current state of inference infrastructure.
-Transparency note: We are being honest about these numbers—they reflect real deployment advantages rather than experimental artifacts.
+While this creates a significant latency difference despite some methods theoretically needing less compute overhead, we consider this a practical advantage reflecting the current state of inference infrastructure and these numbers reflect real deployment advantages of SAEMark.
 Multi-bit scaling.
@@
 The superior information density stems from finer-grained semantic distinctions our framework enables.
-4.4 Adversarial Robustness
+Figure 5: Adversarial robustness. ROC curves showing robust performance against three attack types with varying intensities.
+Figure 6: Multi-bit scaling and information density.
+Watermark acc.
+across different message bits at fixed text length, demonstrating superior information density compared to multi-bit baselines with $\geq$ 90% acc.
+up to 10 bits.
+4.4 Adversarial Robustness and Ablation Studies
+Adversarial Robustness.
 Semantic SAE features provide inherent robustness against paraphrasing attacks.
@@
 Due to space limitations, extended results testing attack intensities up to 50% are provided in Appendix E, demonstrating continued robustness even under stronger attacks.
-4.5 Ablation Studies
-Figure 5: Adversarial robustness. ROC curves showing robust performance against three attack types with varying intensities.
-Figure 6: Multi-bit scaling and information density.
-Watermark accuracy across different message bit lengths at fixed text length, demonstrating superior information density compared to multi-bit baselines with ¿90% accuracy up to 10 bits.
 Figure 7: Framework component ablation studies. Left: Multi-bit watermarking accuracy scaling analysis with ablations. Right: ROC curves for feature concentration hyperparams ($\mu$, $\sigma$).
-Understanding which implementation details drive SAEMark’s empirical success is critically important—our ablation studies validate that these practical optimizations are essential for bridging theoretical bounds and empirical performance, confirming our theoretical predictions about their necessity.
-CheckAlignment filters.
+Ablation Studies primarily involves validating the following components’ contribution to SAEMark’s empirical success: CheckAlignment filters.
 The Range Similarity and Overlap Rate filters prove theoretically grounded and empirically validated.
@@
 5 Conclusion
-SAEMark introduces a fundamental paradigm shift in AI-generated content attribution through feature-based rejection sampling with sparse autoencoder representations.
+SAEMark introduces a paradigm shift in AI-generated content attribution through feature-based rejection sampling.
 Our approach addresses critical limitations of existing watermarking methods by operating entirely through inference-time selection rather than model modification, enabling deployment with API-based services while maintaining superior text quality and detection accuracy.
-Three key advances enable this breakthrough: First, our general framework provides theoretical guarantees that relate watermark success probability to computational budget, independent of the specific feature extractor used.
-Second, the sparse autoencoder instantiation with Feature Concentration Scores captures meaningful semantic patterns that generalize across domains and languages.
-Third, practical optimizations including background feature masking and CheckAlignment filters bridge the gap between theoretical bounds and empirical performance, enabling efficient deployment.
+Three key advances ensures success: First, our general framework provides theoretical guarantees that relate watermark accuracy to computational budget, independent of the specific feature extractor used.
+Second, the sparse autoencoder captures meaningful semantic patterns that generalize across domains and languages and works great as a feature extractor.
+Third, practical optimizations bridge the gap between theoretical bounds and empirical performance, enabling efficient deployment.
 This work establishes that model interpretability tools can be effectively repurposed for content attribution tasks.
@@
 C.1 Single-bit Watermarks
-KGW (Kirchenbauer et al., 2023)
+KGW [16]
 The Key-based Green-list Watermarking (KGW) algorithm is a modern approach for watermarking text generated by LLMs.
-This method builds upon the work of (Kirchenbauer et al., 2023), who introduced a watermarking scheme that divides the token set into ’red’ and ’green’ lists based on a secret key and previously generated tokens.
+This method builds upon the work of [16], who introduced a watermarking scheme that divides the token set into ’red’ and ’green’ lists based on a secret key and previously generated tokens.
 Key features of KGW include the bifurcation of the token set into ’red’ and ’green’ lists, the use of a random seed dependent on a secret key and hash of prior tokens, reweighting of token log-probabilities to favor green tokens, and the introduction of permutation-based reweight strategies.
@@
 The approach offers a balance between watermark embedding and preservation of text quality, addressing challenges faced by previous watermarking methods.
-Unigram (Zhao et al., 2023)
+Unigram [34]
 The Unigram-Watermark and KGW algorithms, both designed for watermarking LLM-generated text, have distinct characteristics.
@@
 This streamlined approach leads to potential efficiency gains in both watermark embedding and detection processes, setting it apart from the more complex K-gram nature of KGW.
-SWEET (Lee et al., 2024)
+SWEET [18]
 The Segment-Wise Entropy-based Embedding Technique (SWEET) is an innovative approach to watermarking code generated by large language models.
@@
 SWEET achieves this by removing low-entropy segments during both the generation and detection of watermarks, effectively balancing the trade-off between detection capability and code quality degradation.
-UPV (Liu et al., 2023)
+UPV [20]
 The key feature of UPV is its use of separate neural networks for watermark generation and detection, addressing the limitation of shared key usage in previous methods.
@@
 This approach ensures high detection accuracy while maintaining computational efficiency, and significantly increases the complexity of forging the watermark, thus enhancing its security in public detection scenarios.
-DIP (Wu et al., 2023)
+DIP [27]
 The Distribution-Preserving Watermarking (DIP) algorithm represents a significant advancement in watermarking techniques for large language models (LLMs).
@@
 Notably, DIP maintains text quality comparable to the original LLM output, owing to its distribution-preserving nature.
-Unbiased (Hu et al., 2023b)
+Unbiased [13]
 Unbiased watermarking and DIP watermarking are closely related concepts in the field of text watermarking for large language models (LLMs).
@@
 This relationship highlights a spectrum of watermarking techniques, where unbiased methods offer a balance between practicality and distribution preservation, while DIP methods provide stronger theoretical guarantees at potentially higher computational costs.
-SynthID (Dathathri et al., 2024)
+SynthID [8]
 SynthID is an advanced watermarking method for large language models (LLMs) that builds upon previous work in generative text watermarking.
@@
 Notably, the method introduces an algorithm to combine generative watermarking with speculative sampling, enabling efficient deployment in high-performance, large-scale production LLMs.
-EXP (Aaronson & Kirchner, 2022)
+EXP [1]
 EXP employs a pseudorandom function $f_{s}()$ with a secret seed $s$ known only to the model provider.
@@
 C.2.1 Baselines
-CTWL (Wang et al., 2023a)
+CTWL [25]
 CTWL is a framework designed to embed multi-bit customizable information into texts produced by large language models (LLMs).
@@
 CTWL seeks to integrate multi-bit information watermarks into LLMs and offers a practical approach for tracing machine-generated texts.
-Waterfall (Lau et al., 2024)
+Waterfall [17]
 Waterfall is a training-free framework designed for robust and scalable text watermarking.
@@
 Waterfall allows for a trade-off between watermark strength and text quality, making it adaptable to various requirements.
-CODEIP (Yoo et al., 2023b)
+CODEIP [30]
 CODEIP embeds a multi‑bit message into generated code by softly biasing token logits during decoding.
@@
 Extraction recovers the ID by re‑hashing and finding the message maximizing cumulative watermark contributions.
-REMARK-LLM (Zhang et al., 2024a)
+REMARK-LLM [32]
 REMARK-LLM introduces a robust watermarking framework for texts generated by large language models.
@@
 The framework incorporates malicious transformations during training to enhance robustness against attacks.
-Provably Robust Multi-bit Watermarking for AI-generated Text (Qu et al., 2024)
+Provably Robust Multi-bit Watermarking for AI-generated Text [22]
 The authors propose a multi‑bit watermarking scheme that embeds a user’s bit‑string ID into LLM‑generated text by first partitioning the message into pseudo‑randomly assigned bit‑segments per token, then biasing “green list” logits seeded by each segment’s value.
@@
 Extraction enumerates only each segment’s possibilities $O(k\cdot 2^{b/k})$, yielding efficient, provably robust watermark recovery under bounded edit distance.
-Robust Multi-bit Text Watermark with LLM-based Paraphrasers (Xu et al., 2024)
+Robust Multi-bit Text Watermark with LLM-based Paraphrasers [28]
 The authors propose a robust multi-bit text watermarking method using LLM-based paraphrasing.
@@
 The decoder acts as a reward model during PPO-based reinforcement learning to fine-tune the encoder, optimizing both the detection of the watermark and the semantic similarity to the original text.
-Robust Multi-bit Natural Language Watermarking through Invariant Features (Yoo et al., 2023a)
+Robust Multi-bit Natural Language Watermarking through Invariant Features [29]
 The authors propose a robust multi-bit natural language watermarking method based on invariant features.
@@
 Appendix D Details of FCS Generation
-This section elaborates on the methodology behind the generation of the Feature Concentration Score (FCS). The process is illustrated in Figure 8, which outlines four key steps.
+This section elaborates on the methodology behind the generation of the Feature Concentration Score (FCS). The process is illustrated in Figure  Figure 8, which outlines four key steps.
 Input: Token sequence $T$, SAE $\theta$ for the entire sequence
@@
 This is achieved through applying the function $argmax_{i}(\phi_{t}\odot m)_{i}$, where $m$ is a mask.
-The output of this function yields the indices corresponding to the most prominent feature, denoted as ”SAE Feature Indices” in Figure  Figure 8.
+The output of this function yields the indices corresponding to the most prominent feature, denoted as "SAE Feature Indices" in Figure  Figure 8.
 Aggregating Most Significant Features
@@
 However, when summarizing the critical features of the entire sequence $T$, redundancies may occur.
-To address this, we employ a set operation to eliminate duplicate entries among the significant features, resulting in a unique collection termed as ”Feature Set SS”.
+To address this, we employ a set operation to eliminate duplicate entries among the significant features, resulting in a unique collection termed as "Feature Set SS".
 Calculating Feature Concentration Score
@@
 Word Deletion Attack
-In the main text, we conducted experiments using the ”maintain content structure” version of the word deletion attack.
+In the main text, we conducted experiments using the "maintain content structure" version of the word deletion attack.
 However, the original word deletion attack involves splitting a paragraph and randomly removing words, which disrupts the structure that watermarking methods rely on, making it harder for the detection system to identify the watermark.
@@
 In our experimental results, we compare two versions of the word deletion attack.
-The ”keep structure” method, represented in a darker color, shows more robust performance with higher AUC values (0.949 at $\epsilon=0.05$ and 0.858 at $\epsilon=0.1$).
-In contrast, the ”not keep structure” method, shown in a lighter color, demonstrates a decline in performance, with AUC values dropping to 0.901 at $\epsilon=0.05$ and 0.825 at $\epsilon=0.1$.
+The "keep structure" method, represented in a darker color, shows more robust performance with higher AUC values (0.949 at $\epsilon=0.05$ and 0.858 at $\epsilon=0.1$).
+In contrast, the "not keep structure" method, shown in a lighter color, demonstrates a decline in performance, with AUC values dropping to 0.901 at $\epsilon=0.05$ and 0.825 at $\epsilon=0.1$.
 These results indicate that preserving the content structure during the attack strengthens the watermark’s resistance, whereas random word deletions that disrupt the structure reduce detection accuracy.
-As shown in the Figure 9, the ”keep structure” method outperforms the ”not keep structure” method in terms of AUC, demonstrating its effectiveness in watermark resistance.
-Figure 9: Word Deletion on SAEMark ROC curves highlighting the performance difference between ”keep structure” and ”not keep structure” methods under word deletion attacks with varying intensities (5%, 10%).
+As shown in the Figure 9, the "keep structure" method outperforms the "not keep structure" method in terms of AUC, demonstrating its effectiveness in watermark resistance.
+Figure 9: Word Deletion on SAEMark ROC curves highlighting the performance difference between "keep structure" and "not keep structure" methods under word deletion attacks with varying intensities (5%, 10%).
 Basic Synonym Substitution Attack
-Our study also examines ”keeping structure” versus ”not keeping structure” approaches in the context of basic synonym substitution attacks, which are less likely to disrupt the content’s structural integrity.
-Figure 10: Basic Synonym Substitution on SAEMark ROC curves comparing ”keep structure” and ”not keep structure” methods under basic synonym substitution attacks at different intensities (5%, 10%).
+Our study also examines "keeping structure" versus "not keeping structure" approaches in the context of basic synonym substitution attacks, which are less likely to disrupt the content’s structural integrity.
+Figure 10: Basic Synonym Substitution on SAEMark ROC curves comparing "keep structure" and "not keep structure" methods under basic synonym substitution attacks at different intensities (5%, 10%).
 Figure 10 shows ROC curves comparing model performance under different conditions, with the original non-structure-preserving method in lighter shades and the modified structure-preserving method in darker hues.
@@
 Due to our algorithm’s prominent performance against context-aware synonym attack. More intensities (20%, 30%, 40%, 50%) are carried upon these kinds of attacking.
-Figure 11: Context-Aware Synonym Substitution on SAEMark ROC curves comparing ”keep structure” and ”not keep structure” methods under basic synonym substitution attacks at different intensities (5%, 10%).
+Figure 11: Context-Aware Synonym Substitution on SAEMark ROC curves comparing "keep structure" and "not keep structure" methods under basic synonym substitution attacks at different intensities (5%, 10%).
 The results of the context-aware watermarking method, shown in  Figure 11 tested under this attack, demonstrate substantial robustness.
@@
 The core research methodology, findings, and analysis remain entirely our own work.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We support all claims with experiment results.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We provide a dedicated limitations section in the Appendix.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We provide proof in methodology section.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We report hyperparameters and we also open-source everything in the linked mentioned in abstract.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We open-source everything in the linked mentioned in abstract.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We report all hyperparameters in the appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification:In "methodology" section, we use t-test as our test method and we report p-values in our experiments.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We report the usage in appendix.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We conform, in every respect, with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We report this in the appendix.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks – we do not train or release any model, and our datasets involved are all from existing, published, peer-reviewed works.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We cite all related assets including models, data, code and baselines.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We provide docs in our repository, including some one-line scripts to reproduce our experiments.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: We describe this in the appendix.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.